# Medformer: A Multi-Granularity Patching Transformer for Medical Time-Series Classification

### v1

@@ Abstract
Medical time series data, such as Electroencephalography (EEG) and Electrocardiography (ECG), play a crucial role in healthcare, such as diagnosing brain and heart diseases. Existing methods for medical time series classification primarily rely on handcrafted biomarkers extraction and CNN-based models, with limited exploration of transformers tailored for medical time series. In this paper, we introduce Medformer, a multi-granularity patching transformer tailored specifically for medical time series classification. Our method incorporates three novel mechanisms to leverage the unique characteristics of medical time series: cross-channel patching to leverage inter-channel correlations, multi-granularity embedding for capturing features at different scales, and two-stage (intra- and inter-granularity) multi-granularity self-attention for learning features and correlations within and among granularities. We conduct extensive experiments on five public datasets under both subject-dependent and challenging subject-independent setups. Results demonstrate Medformer’s superiority over 10 baselines, achieving top averaged ranking across five datasets on all six evaluation metrics. These findings underscore the significant impact of our method on healthcare applications, such as diagnosing Myocardial Infarction, Alzheimer’s, and Parkinson’s disease. We release the source code at https://github.com/DL4mHealth/Medformer.
@@ 1 Introduction
Medical time series refers to sequences of health-related data points recorded at successive times, tracking various physiological signals over time [1, 2]. Classifying medical time series enables continuous monitoring and real-time analysis of a patient’s physiological state, allowing for early detection of abnormalities, accurate diagnosis, timely intervention, and personalized treatment, enhancing patient outcomes and healthcare efficiency [3, 4]. For instance, Electroencephalography (EEG) provides insights into a subject’s neurological status [5, 6], while Electrocardiography (ECG) aids in diagnosing heart conditions [7, 8, 9]. Existing works on medical time series classification mostly utilize handcrafted biomarker extraction [10, 11, 12] and CNN-based models [13, 14, 15, 16], such as EEGNet [17] and TCN [18]. There is a notable absence of effective transformer-based methods for medical time series classification.
@@
Transformers have proven to be effective in time series representation learning across various tasks, including forecasting [19, 20, 21], classification [22, 23], and anomaly detection [24, 25], with a predominant focus on forecasting. While these methods can be applied to medical time series classification, their design motivations and mechanisms may not be ideally suited for this purpose. For instance, as shown in Figure 1, Autoformer [26] and Informer [27] adopt the token embedding method from the vanilla transformer [28], embedding a single cross-channel timestamp as an input token. This approach struggles to capture coarse-grained features along the temporal dimension. Conversely, iTransformer [29] treats the entire series of one channel as an attention token, often overlooking fine-grained temporal features and focusing more on multi-channel correlations. Additionally, PatchTST [30] embeds a sequence of timestamps from one channel as a patch for self-attention, which limits the model’s ability to learn multi-channel correlations.
@@
These existing methods fail to fully exploit the distinctive characteristics of medical time series data, including local temporal dynamics, inter-channel correlations, and multi-scale analysis. Firstly, capturing temporal dynamics requires multi-timestamp input to recognize local temporal patterns with multi-timestamp input, which have been extensively discussed in literature like PatchTST [30]. Secondly, it is essential to harness cross-channel information. For instance, brain activities are generally captured by multi-channel EEG (i.e., 32 channels/electrodes), where each channel monitors a specific brain region. Given that the brain functions as a cohesive unit, the inter-channel correlations (i.e., brain connectome) are crucial for EEG analysis  [31, 32, 33]. Thirdly, representation learning across multiple temporal scales and periods is vital to uncover a broad spectrum of health patterns. Existing research has shown that some signs of disease only exist in specific frequency bands [10, 12].
@@
Figure 1: Token embedding methods. Vanilla transformer, Autoformer, and Informer [28, 26, 27] employ a single cross-channel timestamp as a token; iTransformer [29] utilizes an entire channel as a token; and PatchTST and Crossformer [30, 34] adopt a patch of timestamps from one channel as a token. For medical time series classification, we propose Medformer considering inter-channel dependencies (multi-channel), temporal properties (multi-timestamp), and multifaceted scale of temporal patterns (multi-granularity).
@@
To bridge this gap, we propose Medformer, a multi-granularity patching transformer specifically designed for medical time series classification. Our method introduces three mechanisms to enhance learning ability. Firstly, we propose a novel token embedding approach by implementing cross-channel patching, effectively capturing both multi-timestamp and multi-channel features. Secondly, instead of fixed-length patching, we adopt multi-granularity patching using a list of patch lengths to capture features at different granularities. The multi-granularity approach also simulates different frequency bands and captures band features without relying on handcrafted up/downsampling and band filters. Lastly, we introduce a two-stage (intra- and inter-granularity) multi-granularity self-attention to learn features within specific granularity and the correlations among different granularities. This allows the model to learn features at various granularities and compel them to integrate and complement information.
@@
We conduct extensive experiments using ten baselines and five public datasets, including three EEG datasets and two ECG datasets on detecting Alzheimer’s Disease and cardiovascular diseases, under both subject-dependent and challenging subject-independent setups (Figure 2). The experimental results show that Medformer achieves the best average ranking on all six evaluation metrics across five datasets (Figure 4), demonstrating its superior effectiveness and stability and highlighting its potential for real-world applications.
@@ 2 Related Work
Medical Time Series. Medical time series are a special type of time series data collected from the human body, used for disease diagnosis [3, 7], health monitoring [6, 1], and brain-computer interfaces [2]. Various types of medical time series include EEG [35, 36, 37], ECG [7, 8, 9], EMG [38, 39], and EOG [40, 41]. Each type offers unique capabilities for different healthcare applications. For instance, EEG and ECG can diagnose the health of the brain and heart [35, 7]. Recent research on brain-computer interfaces explores using EEG to control objects, benefiting disabled individuals [2, 42]. Research on medical time series mainly focuses on decoding signals, which involves classifying the hidden information in a sequence of medical time series. Existing works rely on identifying biomarkers and utilizing CNN-based deep-learning models. For example, band features such as relative band power and band correlations [11, 43] have proven useful for EEG-based Alzheimer’s disease diagnosis. Deep learning models like EEGNet [17] and TCN [18, 13] have also demonstrated effectiveness in many medical time series classification tasks.
Table 1: Existing methods do not fully utilize all potential aspects of features in medical time series.
Models Multi- Timestamp Multi- Channel Multi- Granularity Granularity- Interaction
@@
Transformers for Time Series. Existing transformer-based methods for time series can be categorized into two main directions: modifying token embedding methods and self-attention mechanisms, or both. For example, PatchTST [30] uses a sequence of single-channel timestamps as a patch for token embedding. Methods like Autoformer [26], Informer [27], Nonformer [46], and FEDformer [44] develop new self-attention mechanisms or replace the self-attention module to improve learning ability and reduce complexity. Crossformer [34] and iTransformer [29] modify both token embedding methods and self-attention mechanisms. Patching. Patch embedding has been widely used in time series transformers since the proposal of PatchTST [30]. Existing methods of patching, such as Crossformer [34], CARD [21], and MTST [45], inherit from PatchTST [30] and utilize a sequence of single-channel timestamps for patching. This channel-independent patching might benefit learning ability in time series forecasting but may not be as effective in medical time series classification. Multi-Granularity. Existing methods such as Pyraformer [19], MTST [45], Pathformer [47], and Scaleformer [49], utilize multi-granularity embedding to capture features at different scales, allowing models to learn both fine-grained and coarse-grained patterns. We discuss the differences between our method and existing multi-granularity approaches in Appendix F.1.
Medformer includes both novel token embedding and self-attention mechanisms. Figure 1 and Table 1 present a comparison of token embedding methods and feature utilization between our method and existing methods. The components of our method can be easily incorporated into existing methods to improve classification learning ability. For example, cross-channel multi-granularity patching can be integrated with methods that modify self-attention mechanisms, such as Autoformer [26] and Informer [27], for token embedding. Similarly, the two-stage multi-granularity self-attention can be combined with existing multi-granularity methods, like MTST [45], to enhance the learning of inter-granularity features.
@@ 3 Preliminaries and Problem Formulation
Medical time series typically exhibit multiple data levels, including subject, session, trial, and sample levels [13]. In medical time series collected for disease diagnosis tasks, each subject is usually assigned a data label, such as indicating the presence or absence of Alzheimer’s disease. Multiple labels assigned to one subject are also possible if the subject has different diseases. Long sequences of time series(Trial/Session) data from each subject are often segmented into multiple shorter samples for deep learning tasks. Thus, each sample of medical time series generally includes a class label indicating the disease type and a subject ID specifying its origin subject. Given the ultimate goal of diagnosing whether a subject has a particular disease, experimental setups must be meticulously designed to align with real-world medical applications. Diverse experimental setups can yield significantly different results, potentially leading to erroneous conclusions. Here, we introduce two widely employed setups in medical time series classification and highlight their distinctions. Figure 2 provides a simple illustration of these two setups.
@@
Subject-Dependent. In this setup, the division into training, validation, and test sets is based on time series samples. All samples from various subjects are randomly shuffled and then allocated into the respective sets. Consequently, samples with identical subject IDs may be present in the training, validation, and test sets. This scenario potentially introduces “information leakage," wherein the model could inadvertently learn the distribution specific to certain subjects during the training phase. This setup is typically employed for assessing whether a dataset exhibits cross-subject features and has limited applications under real-world medical time-series-based disease diagnosis scenarios. The reason is simple: we cannot know the label of unseen subjects and their corresponding samples during training. Generally, the results of the subject-dependent setup tend to be notably higher than those from the subject-independent setup, often showing the upper limit of a dataset’s learning capability.
@@
Figure 2: Subject-dependent/independent setups (adopted from [13]). In the subject-dependent setup, samples from the same subject can appear in both the training and test sets, causing information leakage. In a subject-independent setup, samples from the same subject are exclusively in either the training or test set, which is more challenging and practically meaningful but less studied.
@@
Subject-Independent. In this setup, the division into training, validation, and test sets is based on subjects. Each subject and their corresponding samples are exclusively distributed into one of the training, validation, or test sets. Consequently, samples with identical subject IDs can only be present in one of these sets. This setup holds significant importance in disease diagnosis tasks as it closely simulates real-world scenarios. It enables us to train a model on subjects with known labels and subsequently evaluate its performance on unseen subjects; in other words, evaluate if a subject has a specific disease. However, this setup poses significant challenges in medical time series classification tasks. Due to the variability in data distribution and the potential presence of unknown noise within each subject’s data, capturing general features across subjects becomes challenging [50, 13, 51, 52]. Even if subjects share the same label, the personal noise inherent in each subject’s data may obscure these common features. Developing a method that effectively captures common features among subjects while disregarding individual noise remains an unsolved problem.
@@
In this work, we evaluate our model mainly in subject-independent setup to better align with real-world applications, and draw the attention of time series researchers to focus on the significant challenges posed by subject-independence. Although this work is not specifically designed for subject-independent problems, our model comprehensively integrates multi-timestamp, multi-channel, and multi-granularity information from each subject. We aim to minimize information loss and enable the model to freely exploit subject-invariant representations. Consequently, our model is well-equipped to tackle the subject-independent challenge to a certain extent, and our results (Section 5) confirm such capability of Medformer.
@@
Next, we present the problem formulation in the context of multivariate medical time series classification for disease diagnosis.
Problem (Medical Time Series Classification).
@@
Consider an input medical time series sample $\mathrm{\bm{x}}_{\text{in}}\in\mathbb{R}^{T\times C}$, where $T$ represents the number of timestamps and $C$ represents the number of channels. We aim to learn a representation $\bm{h}$, which can be used to predict the corresponding label $\bm{y}\in\mathbb{R}^{K}$. Here, $K$ denotes the number of classes that have medical meanings, such as different disease types.
@@ 4 Method
Figure 3: Overview of Medformer. a) Workflow. b) For the input sample $\bm{x}_{\textrm{in}}$, we apply $n$ different patch lengths in parallel to create patched features $\bm{x}_{p}^{(i)}$, where $i$ ranges from 1 to $n$. Each patch length represents a different granularity. These patched features are then linearly transformed into $\bm{x}_{e}^{(i)}$, which are subsequently augmented into $\widetilde{\bm{x}}_{e}^{(i)}$. c) We obtain the final patch embedding $\bm{x}_{(i)}$ by fusing augmented $\widetilde{\bm{x}}_{e}^{(i)}$ with the positional embedding $\bm{W}_{\text{pos}}$ and the granularity embedding $\bm{W}_{\text{gr}}^{(i)}$. Additionally, we design a granularity-specific router $\bm{u}^{(i)}$ to capture integrated information for its respective granularity. We compute both intra-granularity attention, which concentrates within individual granularities, and inter-granularity attention, which leverages the routers to focus across different granularities, for extensive representation learning.
@@
In this section, we first describe the cross-channel multi-granularity patching mechanism for learning spatial-temporal features at different granularities. Next, we analyze the two-stage multi-granularity self-attention mechanism, which leverages features within the same granularity and correlations among different granularities. The architecture of the proposed Medformer is illustrated in Figure 3.
@@
Cross-Channel Multi-Granularity Patch Embedding. From a medical perspective, the brain or heart functions as a cohesive unit, suggesting a naive assumption that there are inherent correlations among different channels in medical time series [31, 32, 33].. Motivated by the above assumption, we reasonably propose multi-channel patching for token embedding, which is different from existing patch embedding methods that embed patches in a channel-independent manner and fail to capture inter-channel correlations [30, 34, 21]. Figure 1 provides an overview comparison of existing token embedding methods and ours. Additionally, existing research on EEG biomarker extraction has shown that certain features are linked to different frequency bands, such as $\alpha,\beta,$ and $\gamma$ bands [10, 12]. This motivates us to embed patch tokens in a multi-granularity way. Instead of using traditional methods like up/downsampling or handcrafted band filtering, multi-granularity patching automatically corresponds to different sampling frequencies, which can simulate different frequency bands and capture band-related features.
@@
Given the above rationales, we propose a novel token embedding approach: cross-channel multi-granularity patching. Given an input multivariate time series sample $\bm{x}_{\text{in}}\in\mathbb{R}^{T\times C}$, and a list of different patch lengths $\left\{{L}_{1},{L}_{2},\ldots,{L}_{n}\right\}$. For the $i$-th patch length $L_{i}$ denoting granularity $i$, we segment the input sample into $N_{i}$ cross-channel non-overlapping patches $\bm{x}_{\text{p}}^{(i)}\in\mathbb{R}^{N{i}\times\left(L_{i}\cdot C\right)}$. Zero padding is applied to ensure that the number of timestamps $T$ is divisible by $L_{i}$, making $N_{i}=\left\lceil{T/L_{i}}\right\rceil$.
@@
A fixed positional embedding $\bm{W}_{\text{pos}}\in\mathbb{R}^{G\times D}$ is generated for positional encoding [28], where $G$ is a very large number. We add $\bm{W}_{\text{pos}}[1:N_{i}]\in\mathbb{R}^{N_{i}\times D}$, the first $N_{i}$ rows of the positional embedding $\bm{W}_{\text{pos}}$, and a learnable granularity embedding $\bm{W}_{\text{gr}}^{(i)}\in\mathbb{R}^{1\times D}$ for the $i$-th patch length $L_{i}$, to obtain the final patch embedding:
@@
where $\bm{x}^{(i)}\in{\mathbb{R}}^{N_{i}\times D}$. Note that the granularity embedding $\bm{W}_{\text{gr}}^{(i)}$ is broadcasted to all $N_{i}$ embeddings during addition.
To reduce time and space complexity, we initialize a router for multi-granularity self-attention (as described later):
@@
Finally, we obtain a list of patch embeddings $\left\{\bm{x}^{(1)},\bm{x}^{(2)},\ldots,\bm{x}^{(n)}\right\}$ and router embeddings $\left\{\bm{u}^{(1)},\bm{u}^{(2)},\ldots,\bm{u}^{(n)}\right\}$ of different patch lengths $\left\{L_{1},L_{2},\ldots,L_{n}\right\}$. We feed the embeddings to the two-stage multi-granularity self-attention.
@@
To reduce the time complexity, we propose a router mechanism and split the self-attention module into two stages: a) intra-granularity and b) inter-granularity self-attention. The intra-granularity stage performs self-attention within the same granularity to capture the distinctive features of each granularity. The inter-granularity stage performs self-attention across different granularities to capture their correlations.
@@
where $\operatorname{Attn}\left(\bm{Q},\bm{K},\bm{V}\right)$ denotes the scaled dot-product attention mechanism in [28]. Note that the router embedding $\bm{u}^{(i)}$ is updated in the same way as the patch embedding $\bm{x}^{(i)}$ to maintain consistency, ensuring that the router can summarize temporal features within the same granularity while the patch embeddings receive global information from the router. The intra-granularity self-attention mechanism enables the model to capture temporal features within the same granularity, facilitating the extraction of local features and correlations among timestamps of the same scale.
Moreover, unlike existing multi-granularity methods such as MTST [45] and Pathformer [47] that use different attention blocks for different granularities, we use a shared attention block for all granularities. This greatly reduces the model’s space complexity and memory consumption, making our method more efficient and scalable.

### v2

@@ Abstract
Medical time series (MedTS) data, such as Electroencephalography (EEG) and Electrocardiography (ECG), play a crucial role in healthcare, such as diagnosing brain and heart diseases. Existing methods for MedTS classification primarily rely on handcrafted biomarkers extraction and CNN-based models, with limited exploration of transformer-based models. In this paper, we introduce Medformer, a multi-granularity patching transformer tailored specifically for MedTS classification. Our method incorporates three novel mechanisms to leverage the unique characteristics of MedTS: cross-channel patching to leverage inter-channel correlations, multi-granularity embedding for capturing features at different scales, and two-stage (intra- and inter-granularity) multi-granularity self-attention for learning features and correlations within and among granularities. We conduct extensive experiments on five public datasets under both subject-dependent and challenging subject-independent setups. Results demonstrate Medformer’s superiority over 10 baselines, achieving top averaged ranking across five datasets on all six evaluation metrics. These findings underscore the significant impact of our method on healthcare applications, such as diagnosing Myocardial Infarction, Alzheimer’s, and Parkinson’s disease. We release the source code at https://github.com/DL4mHealth/Medformer.
@@ 1 Introduction
Medical time series refers to sequences of health-related data points recorded at successive times, tracking various physiological signals over time [1, 2]. Effective classification of MedTS data enables continuous monitoring and real-time analysis of a subject’s physiological state, supporting early abnormality detection, accurate diagnosis, timely intervention, and personalized treatment—ultimately enhancing patient outcomes and healthcare efficiency [3, 4]. For instance, Electroencephalography (EEG) provides insights into a subject’s neurological status [5, 6], while Electrocardiography (ECG) aids in diagnosing heart conditions [7, 8, 9]. Most current MedTS classification approaches rely on handcrafted biomarker extraction [10, 11, 12], convolutional neural networks (CNN)-based models [13, 14, 15, 16], graph convolutional networks (GNNs)-based models[17, 18], or combinations of CNNs and self-attention modules[19, 20]. Notably, effective transformer-based models for MedTS classification remain underexplored.
@@
Transformers have demonstrated strong performance in time series representation learning across tasks such as forecasting [21, 22, 23], classification [24, 25], and anomaly detection [26, 27], with a predominant focus on forecasting. While these methods are applicable to MedTS classification, their design motivations and mechanisms may not fully align with the unique requirements of this domain. For example, as shown in Figure 1, models like Autoformer [28] and Informer [29] adopt the token embedding approach from the vanilla transformer [30], using a single cross-channel timestamp as an input token. This strategy struggles to capture coarse-grained temporal features. In contrast, iTransformer [31] encodes the entire series from one channel as an input token, which can overlook fine-grained temporal features while focusing on multi-channel correlations. Additionally, PatchTST [32] embeds a sequence of timestamps from one channel as a patch for self-attention, limiting the model’s capacity to learn cross-channel relationships.
@@
These existing methods fail to fully exploit the distinctive characteristics of MedTS data, such as local temporal dynamics, inter-channel correlations, and multi-scale feature analysis. First, effective capturing temporal dynamics demands multi-timestamp inputs to capture local temporal patterns, as highlighted in approaches like PatchTST [32] and Crossformer [33]. Second, leveraging cross-channel information is critical; for example, multi-channel EEG data recorded following the International 10–20 system [34] monitors the brain activities, with each electrode/channel corresponding to specific brain regions. Since brain functions are integrated, inter-channel correlations (e.g., brain connectome) are crucial in EEG analysis [35, 36, 37]. Third, representation learning across multiple temporal scales and periods is vital to uncover a broad range of health patterns, as certain disease indicators may only appear within specific frequency bands [10, 12].
@@
Figure 1: Token embedding methods. Vanilla transformer, Autoformer, and Informer [30, 28, 29] employ a single cross-channel timestamp as a token; iTransformer [31] utilizes an entire channel as a token; and PatchTST and Crossformer [32, 33] adopt a patch of timestamps from one channel as a token. For MedTS classification, we propose Medformer considering inter-channel dependencies (cross-channel), temporal properties (multi-timestamp), and multifaceted scale of temporal patterns (multi-granularity).
@@
To bridge this gap, we propose Medformer***We note that this name has been previously used in other domains [38]., a multi-granularity patching transformer designed explicitly for MedTS classification. Our approach introduces three mechanisms to enhance learning capacity. First, we propose a novel token embedding method using cross-channel patching, effectively capturing both multi-timestamp and cross-channel features. To the best of our knowledge, this is the first application of cross-channel patching for transformer embedding in time series analysis. Second, rather than using fixed-length patches, we employ multi-granularity patching with a list of patch lengths, enabling the model to capture features in different scales. This multi-granularity approach could simulate different frequency bands, capturing band-specific features without relying on handcrafted up/downsampling and band filters. Third, We introduce a two-stage (intra- and inter-granularity) self-attention mechanism to capture features within individual granularities and correlations across granularities, enabling complementary information integration across scales.
@@
We conduct extensive experiments using ten baselines across five public datasets, including three EEG datasets and two ECG datasets, focused on detecting diseases such as Alzheimer’s and cardiovascular conditions under both subject-dependent and subject-independent setups (Figure 2). Results show that Medformer achieves the highest average ranking across all six evaluation metrics and five datasets (Figure 4), highlighting its superior effectiveness, stability, and potential for real-world applications.
@@ 2 Related Work
Medical Time Series. Medical time series refers to specialized time series data collected from the human body, commonly used for disease diagnosis [3, 7], health monitoring [6, 1], and brain-computer interfaces (BCIs) [39, 2]. Different MedTS modalities include EEG [40, 41, 42], ECG [7, 8, 9], EMG [43, 44], and EOG [45, 46], each offering distinct capabilities for various medical applications. For example, EEG and ECG data are instrumental in assessing brain and heart health [40, 7]. Recent BCI research explores using EEG to control objects, providing functional support to individuals with disabilities [2, 47]. Unlike general time series research, which predominantly focuses on forecasting tasks [48, 49], MedTS research is centered around signal decoding, which involves classifying hidden information within MedTS sequences. Current approaches often rely on biomarker identification and deep-learning models utilizing CNNs, GNNs, or hybrid models combining CNNs with self-attention modules. For instance, band features such as relative band power and inter-band correlations [11, 50] have proven effective in EEG-based Alzheimer’s disease diagnosis. Deep-learning models like EEGNet [14], EEG-Conformer [20], and TCN [51, 13] have also shown strong performance across various MedTS classification tasks.
Table 1: Existing methods do not fully utilize all potential characteristics in MedTS.
Models Multi- Timestamp Cross- Channel Multi- Granularity Granularity- Interaction
@@
Transformers for Time Series. Transformer has demonstrated its strong learning and scaling-up ability in many domains, including natural language processing [30, 57] and computer vision [58, 59]. Existing transformer-based methods for time series analysis can be categorized into two main directions: modifying token embedding methods and self-attention mechanisms, or both. For example, PatchTST [32] uses a sequence of single-channel timestamps as a patch for token embedding. Methods like Autoformer [28], Informer [29], Nonformer [54], and FEDformer [52] develop new self-attention mechanisms or replace the self-attention module to improve learning ability and reduce complexity. Crossformer [33] and iTransformer [31] modify both token embedding methods and self-attention mechanisms. Patching. Patch embedding has been widely used in time series transformers since the proposal of PatchTST [32]. Existing methods of patching, such as Crossformer [33], CARD [23], and MTST [53], inherit from PatchTST [32] and utilize a sequence of single-channel timestamps for patching. This channel-independent patching might benefit learning ability in time series forecasting but may not be as effective in MedTS classification. Multi-Granularity. Existing methods such as Pyraformer [21], MTST [53], Pathformer [55], and Scaleformer [60], utilize multi-granularity embedding to capture features at different scales, allowing models to learn both fine-grained and coarse-grained patterns. We discuss the differences between our method and existing multi-granularity approaches in Appendix G.1.
Medformer includes both novel token embedding and self-attention mechanisms. Figure 1 and Table 1 present a comparison of token embedding methods and feature utilization between our method and existing methods. The components of our method can be easily incorporated into existing methods to improve classification learning ability. For example, cross-channel multi-granularity patching can be integrated with methods that modify self-attention mechanisms, such as Autoformer [28] and Informer [29], for token embedding. Similarly, the two-stage multi-granularity self-attention can be combined with existing multi-granularity methods, like MTST [53] and Pathformer [55], to enhance the learning of inter-granularity features.
@@ 3 Preliminaries and Problem Formulation
Disease Diagnosis with MedTS. Medical time series data typically exhibit multiple hierarchical levels, including subject, session, trial, and sample levels [13]. In disease diagnosis tasks using MedTS, each subject is usually assigned a single label, such as indicating the presence or absence of Alzheimer’s disease. However, multi-labeling may be necessary when a subject has co-occurring conditions [61]. Notably, a subject’s medical or physiological state remains relatively stable over time (or within short periods without significant change). For instance, a subject diagnosed with Alzheimer’s disease is expected to retain that diagnosis for many years. If the subject also has Parkinson’s disease, a multi-label learning approach is required, which essentially conducts classification tasks for each disease independently if they are not mutually exclusive.
@@
Since long sequences of time series data (e.g., trials or sessions) are often segmented into shorter samples for deep learning training, all samples from a single subject should ideally retain the same medical condition label. Thus, each MedTS sample typically includes a class label indicating a specific disease type and a subject ID indicating its original subject. Given the ultimate goal of diagnosing whether a subject has a particular disease, experimental setups must be carefully designed to reflect real-world clinical applications. Diverse setups can yield markedly different outcomes, potentially leading to misleading conclusions. Here, we introduce two widely used setups in MedTS classification and clarify their distinctions. Figure 2 provides a simple illustration of these two setups.
@@
Subject-Dependent. In this setup, the division into training, validation, and test sets is based on time series samples. All samples from various subjects are randomly shuffled and then allocated into the respective sets. Consequently, samples with identical subject IDs may be present in the training, validation, and test sets. This scenario potentially introduces “information leakage," wherein the model could inadvertently learn the distribution specific to certain subjects during the training phase. This setup is typically employed to assess whether a dataset exhibits cross-subject features and has limited applications under real-world MedTS-based disease diagnosis scenarios. The reason is simple: we cannot know the label of unseen subjects and their corresponding samples during training. Generally, the results of the subject-dependent setup tend to be notably higher than those from the subject-independent setup, often showing the upper limit of a dataset’s learning capability.
@@
Figure 2: Subject-dependent/independent setups (figure adopted from our previous work [13]). In the subject-dependent setup, samples from the same subject can appear in both the training and test sets, causing information leakage. In a subject-independent setup, samples from the same subject are exclusively in either the training or test set, which is more challenging and practically meaningful but less studied.
@@
Subject-Independent. In this setup, the division into training, validation, and test sets is based on subjects. Each subject and their corresponding samples are exclusively distributed into one of the training, validation, or test sets. Consequently, samples with identical subject IDs can only be present in one of these sets. This setup holds significant importance in disease diagnosis tasks as it closely simulates real-world scenarios. It enables us to train a model on subjects with known labels and subsequently evaluate its performance on unseen subjects; in other words, evaluate if a subject has a specific disease. However, this setup poses significant challenges in MedTS classification tasks. Due to the variability in data distribution and the potential presence of unknown noise within each subject’s data, capturing general task-related features across subjects becomes challenging [62, 13, 63, 64]. Even if subjects share the same label, the personal noise inherent in each subject’s data may obscure these common features. Developing a method that effectively captures common features among subjects while disregarding individual noise remains an unsolved problem.
@@
In this work, we evaluate our model mainly in the subject-independent setup to better align with real-world applications, aiming to draw attention within the time series research community to the substantial challenges posed by this setup. While our model is not specifically tailored to address subject-independent problems, it integrates multi-timestamp, cross-channel, and multi-granularity features in MedTS, enhancing its capacity to capture subject-invariant representations. Consequently, our model is well-equipped to tackle the subject-independent challenge to a certain extent, and our results (Section 5) confirm such capability of Medformer.
@@
We next present the problem formulation for multivariate medical time series classification in the context of disease diagnosis.
Problem (MedTS Classification).
@@
Consider an input MedTS sample $\bm{x}_{\text{in}}\in\mathbb{R}^{T\times C}$, where $T$ denotes the number of timestamps and $C$ represents the number of channels. Our objective is to learn an encoder that can generate a representation $\bm{h}$, which can be used to predict the corresponding label $\bm{y}\in\mathbb{R}^{K}$ of the input sample. Here, $K$ denotes the number of medically relevant classes, such as various disease types or different stages of one disease.
@@ 4 Method
Figure 3: Overview of Medformer. a) Workflow. b) For the input sample $\bm{x}_{\textrm{in}}$, we apply $n$ distinct patch lengths in parallel to create patched features $\bm{x}_{p}^{(i)}$, where $i$ ranges from 1 to $n$. Each patch length represents a unique granularity. These patches are then projected into $\bm{x}_{e}^{(i)}$ and subsequently augmented to form $\widetilde{\bm{x}}_{e}^{(i)}$. c) We obtain the final embedding $\bm{x}^{(i)}$ by combining the augmented $\widetilde{\bm{x}}_{e}^{(i)}$ with both the positional embedding $\bm{W}_{\text{pos}}$ and the granularity embedding $\bm{W}_{\text{gr}}^{(i)}$. Additionally, a granularity-specific router $\bm{u}^{(i)}$ is designed to capture integrated information for each respective granularity. We then perform intra-granularity self-attention, focusing on individual granularities, and inter-granularity self-attention, using the routers to facilitate communication across different granularities.
@@
In this section, we first describe the cross-channel multi-granularity patching mechanism for learning spatial-temporal features in different scales. Next, we analyze the two-stage multi-granularity self-attention mechanism, which leverages features within the same granularity and correlations among different granularities. The architecture of the proposed Medformer is illustrated in Figure 3.
@@
Cross-Channel Multi-Granularity Patch Embedding. From the medical perspective, the brain or heart functions as a cohesive unit, suggesting a naive assumption that there are inherent correlations among different channels in MedTS [35, 36, 37], as each channel represents the activities of distinct brain or heart regions. Motivated by the above assumption, we reasonably propose multi-channel patching for token embedding, which is different from existing patch embedding methods that embed patches in a channel-independent manner and fail to capture inter-channel correlations [32, 33, 23]. Figure 1 provides an overview comparison of existing token embedding methods and ours. Additionally, existing research on EEG biomarker extraction has shown that certain features are linked to different frequency bands, such as $\alpha,\beta,$ and $\gamma$ bands [10, 12]. This motivates us to embed patch tokens in a multi-granularity way. Instead of using traditional methods like up/downsampling or handcrafted band filtering, multi-granularity patching automatically corresponds to different sampling frequencies, which can simulate different frequency bands and capture band-related features.
@@
Given the above rationales, we propose a novel token embedding approach: cross-channel multi-granularity patching. Given an input multivariate MedTS sample $\bm{x}_{\text{in}}\in\mathbb{R}^{T\times C}$, and a list of different patch lengths $\left\{{L}_{1},{L}_{2},\ldots,{L}_{n}\right\}$. For the $i$-th patch length $L_{i}$ denoting granularity $i$, we segment the input sample into $N_{i}$ cross-channel non-overlapping patches $\bm{x}_{\text{p}}^{(i)}\in\mathbb{R}^{N_{i}\times\left(L_{i}\cdot C\right)}$. Zero padding is applied to ensure that the number of timestamps $T$ is divisible by $L_{i}$, making $N_{i}=\left\lceil{T/L_{i}}\right\rceil$.
@@
A fixed positional embedding $\bm{W}_{\text{pos}}\in\mathbb{R}^{G\times D}$ is generated for positional encoding [30], where $G$ is a very large number. We add $\bm{W}_{\text{pos}}[1:N_{i}]\in\mathbb{R}^{N_{i}\times D}$, the first $N_{i}$ rows of the positional embedding $\bm{W}_{\text{pos}}$, and a learnable granularity embedding $\bm{W}_{\text{gr}}^{(i)}\in\mathbb{R}^{1\times D}$, to obtain the final patch embedding for the $i$-th granularity with patch length $L_{i}$:
@@
where $\bm{x}^{(i)}\in{\mathbb{R}}^{N_{i}\times D}$. Note that the granularity embedding $\bm{W}_{\text{gr}}^{(i)}$ aims to distinguish among granularities and is broadcasted to all $N_{i}$ embedding rows with $D$ dimension during addition.
To reduce time and space complexity, we initialize a router to be used in the multi-granularity self-attention (as described later) for each granularity:
@@
Finally, we obtain a list of final patch embeddings $\left\{\bm{x}^{(1)},\bm{x}^{(2)},\ldots,\bm{x}^{(n)}\right\}$ and router embeddings $\left\{\bm{u}^{(1)},\bm{u}^{(2)},\ldots,\bm{u}^{(n)}\right\}$ for different granularities with patch lengths $\left\{L_{1},L_{2},\ldots,L_{n}\right\}$. We feed the embeddings to the two-stage multi-granularity self-attention.
@@
To reduce the time complexity, we propose a router mechanism and split the self-attention module into two stages: a) intra-granularity self-attention and b) inter-granularity self-attention. The intra-granularity stage performs self-attention within the same granularity to capture the distinctive features of each granularity. The inter-granularity stage performs self-attention across different granularities to capture their correlations.
@@
where $\operatorname{Attn}\left(\bm{Q},\bm{K},\bm{V}\right)$ denotes the scaled dot-product self-attention mechanism in [30]. Note that the router embedding $\bm{u}^{(i)}$ is updated concurrently with the patch embedding $\bm{x}^{(i)}$ to maintain consistency, ensuring that the router effectively summarizes each granularity’s features in the current training step and is ready for the subsequent inter-granularity self-attention. The intra-granularity self-attention mechanism allows the model to capture temporal features within a single granularity, facilitating the extraction of local features and correlations among timestamps at the same scale.